x + 1/2 = 3x + 5/6
Answer by checkley77(12844)

x + 1/2 = 3x + 5/6
X-3X=5/6-1/2
-2X=5/6-3/6
-2X=2/6
-2X=1/3
X=(1/3)/-2
X=1/3*-1/2
X=-1/6 ANS.
PROOF:
-1/6+1/2=3*-1/6+5/6
-1/6+3/6=-3/6+5/6
2/6=2/6
1/3=1/3